OpenStudy (anonymous):

Solve the equation for x. If necessary, round your answer to two decimal places. Do not include "x =" in your answer. (1/5)^x = 1/25

OpenStudy (anonymous):

You can solve this easily with knowledge that 25 is 5^2, but using logs: \[(\frac{1}{5})^x=\frac{1}{25}\] \[x \log \frac{1}{5}=\log\frac{1}{25}\] \[x=2\]

OpenStudy (anonymous):

(1/5)^x=1/25 (1/5)^x=(1/5)^2 x=2

OpenStudy (anonymous):

5x5 = 25 ==> 5^2